OpenStudy (jhuckaby):

Mary is buying several items that cost $128.25 total. She is using a store coupon for 35% off her purchases. She has to pay 4% sales tax. Calculate the total cost of the items.

OpenStudy (holsteremission):

\[\mathrm{total}=\mathrm{subtotal}\times(1-\mathrm{discount})\times(1+\mathrm{tax})\]Why? Discounts are taken into account before tax. So if the subtotal is \(x\) and the discount is \(d\), then you remove \(dx\) from \(x\), i.e. \(x-dx=x(1-d)=x^*\), where \(x^*\) denotes the new subtotal. The sales tax \(s\) then kicks in, adding \(s\) times the value of the subtotal, i.e. \(x^*+sx^*=x^*(1+s)=x(1-d)(1+s)\).

OpenStudy (mathstudent55):

The items cost $128.25 before the discount and before the tax. To find the discounted price, first, subtract 35% from 100% to get 65%. 65% is the same as 0.65 as a decimal. Then multiply the cost before discount and taxes by 0.65. To add the tax, first add 100% and 4% to get 104%. 104% is the same as 1.04 as a decimal. Then multiply the discounted price by 1.04.